OpenStudy (anonymous):

What is the approximate circumference of a pizza pan with a diameter of 13.2 inches? (Use π = 3.14) A. 20.724 inches B. 82.896 inches C. 62.172 inches D. 41.448 inches

OpenStudy (anonymous):

@Directrix

OpenStudy (triciaal):

circumference = 2 * pi* r r = 1/2 *diameter

OpenStudy (anonymous):

1.6=r

OpenStudy (triciaal):